OPEN JULY 4TH
11AM to 6PM

Grinkers Grand Palace
Friday, December 5th 2014
Star Wars cockpit
Partial monitor collapse
Replaced 2N3792 transistor
Star Wars cockpit is 100% operational

Liberator
Game would not boot
Replaced main 7amp fuse
Replace power cord
Liberator is 100% operational

Rip Off
Partial monitor collapse
Replaced 2N3792 transistor
Rip Off is 100% operational

Blue Shark
Screen was going out intermittently and sound was not correct
Replaced obsolete connector on main board
Removed pot from speaker line
Blue Shark is 100% operational

Space Duel
Game was playing blind
Replaced two 2N3716 and three 2N3792 transistors
Replaced two blown fuses
Replaced two capacitors on the high voltage board
Space Duel is 100% operational

Black Hole
Replaced motor on back glass spinner

Comet
Replaced burned connector and pin on power supply board

Gap in updates = too much time fixin', not enough time logging. We'll try to do better.

Saturday February 1st, 2014
Frogger cocktail
Player Two joystick not registering backward movement.
Joystick microswitch replaced.
Frogger cocktail is 100% operational.

Race Drivin'
Received reports of shifter not functioning correctly.
Calibrated shifter - all appears to be functioning correctly.
Rave Drivin' is 100% operational.

Mousetrap
We are preparing mousetrap for its first release to the arcade.
PCB that came with the game was fried.
Purchased and replaced pcb.
Game played but monitor in bad shape.
Rebuilt k4900 chassis.
Monitor looking good but now noticed grapgical glitches.
Reseated all roms, resoldered header pins on main pcb.
Graphics glitches gone and game now plays 100%.
Status: Moustrap needs coin mech installed then will be ready for the arcade.

Operation Wolf
We are preparing Operation Wolf for its first release to the arcade.
Gun assembly loose and wobbly. Tightened assembly.
Adjusted +5 power supply.
Game boots up with ROM error.
Reseaeted roms but ROM error remains.
Pulled and tested all ROMS. ROM data verified successfully.
Replace ROMS.
Game still boots up with ROM error.
Status: Continued effort needed.

Earthshaker
Left flipper not working.
Replaced lower left flipper coil.
Earthshaker is 100% operational.

Terminator 2: Judgement Day
Back box lighting connertor faulty causing general illuminate issues.
Replaced jack and plug.
Terminator 2: Judgement Day is 100% operational.

Tron
Backlight showing MCP burned out.
Replaced fluorescent bulb.
Tron is 100% operational.

Heavy Barrel
Marquee bulb burned out.
Replaced fluorescent bulb.
Heavy Barrel is 100% operational.

Sinistar
Marquee bulb burned out.
Replaced fluorescent bulb.
Sinistar is 100% operational.

Paperboy
Marquee bulb burned out.
Replaced fluorescent bulb.
Paperboy is 100% operational.

t2 - adjust gun switch - reasearch more.

Tetris
Measured for new plexi
Other than missing front plexi, Tetris is 100% operational.

Anti-Aircraft
Measure for replacement lower backdoor.
Other than missing lower backdoor, Tetris is 100% operational.

Xenon
Replaced approximately twelve burned out bulbs.
One bulb not lighting, need to identify circuit.
Other than the one bulb, Xenon is 100% operational.

Pinball Dolly
Modified our pinball dolly handles to pivot enabling easier moving of pinball games.


Saturday January 18th, 2014
Circus Charlie
Game would not power up.
Identified faulty switching power supply and replaced.
Status: Circus Charlie is 100% operational.

Thief
Game would power up but no image on monitor.
Identified blown fuse on power supply and replaced.
Status: Thief is 100% operational.

Frogger - upright
Various color problems rendering game incorrect and difficult to play.
Color adjustments and various other fix attenpts yielded no improvement.
Located, purchased and installed replacement PCB.
Status: Frogger is 100% operational. All colors are now correct.

Cheyenne
Mine level was not displaying correctly.
Pulled PCB stack and reseated all ROMS.
Mine level now appeared correct however gun shots now wouldn't register.
Calibrated gun.
Status: Cheyenne is 100% operational.

Gravitar
Graphics on monitor would flicker significantly during in-game explosions or when cabinet was otherwise vibrated.
Fixed cold solder joints on deflection board.
Status: Gravitar is 100% operational.

Firetruck
Very low volume. Rear half of fire truck steeting would steer in only one direction.
Replaced volume pot and replaced amplifier.
Cleaned steering optics.
Status: Firetruck is 100% operational.

1942
Right side of control panel was loose.
Resecured control panel bolts.
Status: 1942 is 100% operational.

Star Wars pinball
Right side flipper would stick in open position intermittently.
Replaced right side flipper coil stop and armature.
Need to purchase replacement SC board.
Status: Star Wars pinball is presently offline.

Added porkchop to Astro Blaster right side coin mech

Thursday, June 6th
High Speed
Reconnected knocker.
Voltage present on ramps. Identified issue due to pinched flasher wire under right hideout ramp. Unpinched wire.
Replaced lamp in center jet bumper.
Tweaked exit ramp and secured two bumpers on the playfield that had come loose. Shots now go up ramp like they should.
One of the right hideout switches is not correct for that application and should be replaced. Need to order replacement. Adjusted existing switch and is working fine at present.

Comet
Replaced all jet bumper and two playfield lamps under the ramp with LED lamps.
Adjusted jet bumper sockets for more secure connection.
Dressed the flipper contacts with contact file.
Tightened several playfield bumpers that were loose.

Pinbot
Adjusted the vortex and back ball guide so the ball properly goes up the vortex ramp when launched.
Replaced a general illumination lamp in back left corner.

Terminator
Replaced F102 right flipper fuse.
Replaced back EMF diode across the right flipper coil.

Saturday, June 1st
Super Sprint
Blue pedal releases but still reads as depressed.
Cleaned dirt and debris from pedal mech.
Adjusted pot.
Status: Super Sprint pedal is working correctly. Game is 100% operational.

Phoenix cocktail
Scrambled screen and noise from speaker.
Adjusted verticle hold on monitor.
Fixed broken speaker wire.
Status: Phoenix cocktail is 100% operational.

Star Wars
Loud fan, likely close to failure.
Replaced fan.
Status: Star Wars is 100% operational.

Robotron
Left joystick grinding at top left position.
Adjusted joystick.
Status: Robortron is 100% operational.

Asteroids Deluxe
Sound cracking.
Status: May need an AR board rebuild. Aded to fix list.

Tron
Garbled sprites in MCP level at MCP blocks.
Status: Possible bad object RAM. Monitor and consider replacing or reseating RAM.

MC Cockpit
Trackball clicking slightly on verticle movements.
Found slight imperfection on roller shaft. It is a new shaft with little wear.
Status: Shaft will likely smooth out over time. Monitor.

Gravitar
Fuzzy vector images.
Replaced missing High Voltage cage cover.
Status: Gravitar is 100% operational.

Asteroids
Vectors images squished at left and right edges of monitor.
Adjusted linearity on deflection board.
Status: Adjustment helped some but did not resolve. Need to replace linearity pots on deflection board.

Heavy Barrel
Screen very dark.
Tested tube, all guns weak.
Rejuved monitor. Note that this is second rejuve. Monitor was first rejuved around Grinker's opening day. Monitor likely near end of life.
Status: Heavy Barrel is 100%. Monitor for need for additional rejuves and potential monitor replacement.

Skull and Crossbones
Left joystick not operable.
Broken actuator on joystick.
Fixed MacGyver style with electrical tape.
Status: Skull and Crossbones is functional. Monitor for future proper repair.

1942
Control panel loose. Volume too low.
Tightened screws to secure control panel.
Increased volume.
Status: 1942 is 100% operational.

High Speed pinball
Replaced main power fuse with correct fuse.
Flipper power fuse blew due to upper right flipper plunger link broken and not activating end of stroke switch.
Flipper coil in stock is used. Need to replace with new.

Terminator pinball
Replaced F109.
Needs connector and header replaced on power driver board

Saturday, May 25th
Race Drivin'
Game playable but monitor dull and colors not correct.
Replaced bad capacitors at monitor.
Reflowed solder where needed on monitor PCB.
Rejuvinated tube.
Adjusted colors.
Status: Race Drivin' monitor looks great. Need to order rubber boot for shifter. Need to install NVRAM and 10K Pot.

Tempest
Game not booting.
Found four dead transistors on monitor cage and a blown f600 fuse.
Replaced transistors and fuse. Game boots but only 1/2 monitor displays.
Status: Tempest is offline. Further investigation required

Sunday, May 19th
Q*Bert
Random lockups / power issues.
Bypassed filter board by splicing harness wires.
Ran Q*Bert for 24 hours. No problems.
Status: Q*Bert is 100% operational.

Satan's Hollow
Voltage issue at PCB.
Status: Satan's Hollow is down. Investigate suitcase power brick.

Discovered I was hungry and ate a yummay sammich (#10).
Problem solved, no longer hungry.

Saturday, May 18th
Race Drivin'
Power cord missing groung plug.
Replaced power cord plug end.
Tested voltage: 5.09 at PCB.
Game boots and plays.
Status: Playable but needs: (a) zero-power RAM chips to save calibration settings, (b) monitor recapped/adjusted, (c) rubber boot for shifter.

Space Invaders
Screen scrambled.
Adjusted horizontal hold.
Status: Space Invaders is 100% operational. Replace horizontal hold pot if problem reoccurs.

Super Mario Brothers
Replaced broken player 1 start button.
Status: Super Mario Brothers is 100% operational.

Q*Bert
Random lockups / power issues.
Inspected filter board. Identified bad resistor.
Status: Q*Bert is down. Need to bypass filter board by splicing harness wires.

Tapper
Game booting to cross hatch pattern.
Disconnected reset wire - no change.
Changed 100,000 MF and 56,000 MF filter caps.
Game booted successfully.
Controls unresponsive.
Found and resoldered broken ground wire on control panel.
Controls now working.
Status: Tapper is 100% operational. Need to replace red ball joystick with NOS tapper handle joystick.

Gravitar
Game non-functional.
Swapped Space Dual PCB into Gravitar.
Worked 100%.
Issue appears to be with Gravitar PCB.
Status: Gravitar is down. Need to research Gravitar PCB fixes.

Saturday, May 11th
Paperboy
Blurry and dim screen.
Tested tube with rejuvenator.
Red, green and blue testing bad.
Used rejuvenator to get all three colors testing good.
Status: Paperboy is 100% operational.

Gravitar
Game is playing dark, no picture.
Found no HV after testing B+.
Pulled HV cage, tested flyback with ring tester, found no bad components in HV cage.
Found out of spec resistor on deflection board, replaced. Spot killer on.
Tested Space Duel board, works fine, no spot killer. Reseated all edge connectors.
Status: Possible board problem. Continue efforts next Saturday.

Tapper
Booting to crosshatch pattern.
Replaced power cord (missing ground plug and insulation in a couple areas).
Found test switch not wired up, fixed.
Rebuilt 90412 power supply board.
Tested voltages on pcb, adjusted to 5.1.
Status: Still booting to cross hatch pattern. Further investigation needed.

Vanguard
Replaced marquee starter and bulb.
Status: Vanguard is 100% operational.

Saturday, May 4th
Gravitar
Scrambled vector graphics on bottom portion of screen.
Replaced bad diode and bad transistor on k6100 deflection board.
Status: Gravitar is 100% operational.

Qbert
Intermittent problems with game crashing to garbled screen.
Confirmed that voltages to board are good.
Reseated all edge connectors.
Status: Qbert is operational. Monitor behavior to determine if problem persists.

Sprint 2
Screeching sound and scrambled garbage on screen.
Identified only ~3volts on PCB.
Suspect problem with Lm323k voltage regulator.
Status: Sprint 2 is down. Lm323k voltage regulator has been ordered.

Elevator Action
Character sprites were not displaying correctly.
Purchased and installed replacement PCB.
New PCB fixed sprite issue, but then we identified that colors were not correct; no blue.
Checked tube with rejuvenator. All color guns functioning.
Hooked pattern generator to chassis and same results; no blue.
Pulled monitor pcb and checked usual suspects (color transistors, color pot). All checked good.
Found cold solder on blue pin of neck socket. Resoldered and adjusted color pots.
Status: Elevator Action is 100% operational.

Alpine Ski
Replaced power cord.
Status: Alpine Ski is 100% operational.

Ice Cold Beer
Not registering ball drops correctly in hole #3.
Adjusted swiches and confirmed correct operation.
Status: Ice Colde Beer is operational. Monitor behavior to determine if problem persists.

Donkey Kong cocktail
Nintendo coin mech not working correctly.
Replaced with alternate coin mech from backup coin door.
Status: Donkey Kong cocktail is 100% operational.

Ice Cold Beer
Not registering ball drops correctly in hole #3.
Adjusted swiches and confirmed correct operation.
Status: Ice Cold Beer is operational. Monitor behavior to determine if problem persists.

April
We did lots of stuff in April but we failed to log our efforts.
We'll do better in May.

Saturday, March 30th
Krull
Game wouldn't power up.
Replaced power cord end.
Game powered up but with no sound.
Reseated edge connectors, sound returned.
Status: Krull is 100% operational.

Q*bert
Monitor garbled.
Adjusted 5V on power supply.
Status: Q*bert is 100% operational.

Missile Command
Color problem affecting text at bottom and playability of some levels.
Adjusted color levels.
Status: Missile Command is 100% operational.

Thursday, March 28th
Q*bert
Bottom half of joystick disconnected from control panel. Game not playable.
Replaced and tighened nuts on joystick bolts.
Status: Q*bert is 100% operational.

High Speed
Replaced broken right flipper rubber.
Status: High Speed is 100% operational.

Frenzy
PCB had disconnected from support railing.
Removed PCB and reconnected to railing.
Game booted to white screen.
Reseated edge connectors and molex connector.
Status: Frenzy is 100% operational.

Friday, March 22nd
T2 Pinball
Shooter loose causing balls to launch very weakly.
Used toothpick and wood glue trick to resecure three screws at launcher solenoid platform.
Status: T2 Pinball is 100% operational.

Pinbot
Two balls were being set in launcher with each new ball.
Ball from tilt mechanism had been put in play. Returned ball to tilt mechanism.
Replaced a burned out playfiled bulb.
Bank of switches not working. Resoldered disconnected wire.
Problems with general illumination operation. Resoldered loose wire.
Non functional switch. Resoldered connection.
Status: Pinbot is 100% operational.

Xenon
Replaced burned out playfield bulb.
Adjusted credits per coin. Was giving ten credits per coin. Adjusted to one credit per coin.
Status: Xenon is 100% operational.

Q*bert
Control panel not attached. Screws had loosened and disconnected.
Added lock washers. Resecured control panel.
Status: Q*bert is 100% operational.

Sunday, March 17th
Super Off Road
Replaced molex connector on wiring for Nitro buttons.
Placed and secured control panel and gas pedal components.
Status: Super Off Road joined the arcade for the first time on March 17th, 2013. Super Off Road is 100% operational.

Saturday, March 16th
Pac-Man
Monitor non-functional. Reset anode cup off that had disconnected from monitor tube.
Monitor not diplaying red. Rejuvinated red gun.
Status: Pac-Man is 100% operational.

Ms. Pac-Man
Monitor displaying test screen only. Reconnected VRAM address daughter card that had disconnected from main board.
Status: Ms. Pac-Man is 100% operational.

Xevious
Joystick "up" direction not working. Replaced broken leaf switch.
Status: Xevious is 100% operational.

Donkey Kong cocktail
Monitor scrolling vertically. Replaced broken v-hold pot.
Loud sound distortion and "scratching". Resoldered broken speaker wire to transformer.
Status: Donkey Kong cocktail is 100% operational.

Airwolf
Monitor scrolling vertically. Adjusted vertical hold to stabalize picture.
Status: Airwolf is 100% operational.

Super Offroad
Replaced bad power supply.
Replaced burnt power connector.
Replaced marquee bulb.
Status: Game works with exception of Nitro buttons. Needs 4 position molex connector that was cut off.

Ten Yard Fight
Replaced power cord that had exposed wires.
Status: Power supply has failed. Requires replacement. Will order parts week of March 17th.

Baby Pac-Man
Replaced blown 1 amp fuse under playfield.
Status: Intermittent problems continue with monitor and ball launching from launching solenoid. Expect MPU board is bad. Will locate and order parts week of March 17th

Saturday, March 9th
Q*bert
Replaced power cord.
Set dip switches to quarters instead of free play.
Changed marquee bulb.
Resecured control panel.
Status: Qbert joined the arcade for the first time on March 9th, 2013. Q*bert is 100% operational.

Pac-Man
Replaced microswitched joystick with vintage leaf joystick.
Swapped older control panel overlay with cleaner control panel overlay.
Status: Pac-Man is 100% operational.

Qix
Game would not boot. Replaced faulty 6514 Ram chips.
Status: Qix is 100% operational.

Frogger cocktail
Graphics were jumbled. Replaced faulty power supply.
Status: Frogger is 100% operational.

Saturday, March 2nd
Sprint 2
Sprint non-functional since purchase.
Replaced LM323K voltage regulator.
Cleaned and rebalanced monitor tube.
Rewired left side gear shifter.
Swapped right side steering wheel.
Swapped right side steering board.
Replaced right side gear shift bezel.
Status: Sprint 2 joined the arcade for the first time on March 2nd, 2013. Sprint 2 is 100% operational.

Tempest
Tempest was non-functional. Loud hum coming from speaker.
Replaced Big Blue capacitor.
Replaced missing cone on two player button.
Status: Tempest is 100% operational.

Moon Patrol
Left jump button was sticking.
Button had been hacked together. Replaced button.
Status: Moon Patrol is 100% operational.

Battlezone (A)
Game was playing dark (monitor not displaying).
Replaced 3716 and 3792 deflection transistors on the monitor.
Bypassed transistor holder on monitor frame.
Fixed loose pin on monitor molex connector.
Status: Battlezone is 100% operational.

Tron
Marquee light not functioning.
Replaced lamp starter.
Status: Tron marquee light is now functioning. Tron is 100% operational.

Thursday, February 28th
T2 pinball
Multiball not working correctly.
Replaced fried chip.
Status: T2 is now 100% operational

Xenon pinball
Right slingshot not operational.
Resoldered disconnected wire at right slingshot.
Status: Xenon is 100% operational

Wednesday, February 27th
Xenon pinball
Several lights not working on playfield.
Replaced lamp board. All lights now working correctly.
Status: Xenon is 100% operational